Steven has an average of 76% on his first 5 tests. What is the lowest score that he can get on his next test to have an average of 80% on all of his tests?
a. 76% b. 80% c. 90% d. 100%

Answer by checkley77(12569)

(76*5+X)/6=80
(380+X)/6=80
380+X=6*80
380+X=480
X=480-380
X=100 IS THE NECESSARY GRADE FOR THE NEXT TEST.